Bwindi Impenetrable Forest (10.01.2013 to 12.31.2013)

You are here

PDF versionPDF version
Year: 
2013
Quarter: 
October - December
Communications activities: 

No communication events within this quater.

Unusual Events at the TEAM Site: 

No unusual events within this quater.

New Species: 

No new species within this quarter.

Protocol Activities: 

Vegetation monitoring:

Measurement error estimation for the diameter takers was conducted. A refresher training for all the crew was also conducted before plot re-measurement. Three (5, 3 and 4) of the six vegetation plots were re-measured. Data for two plots has been entered and yet to be uploaded on to the TEAM portal.

 

Climate monitoring:

The climate station was repaired and a few additions were made. This was done in liaison with the TEAM technical personnel. The battery house was modified to allow improved aeration (crucial to ensure battery long life). A new data logger and solar regulator were installed. A new HC2-S3 temperature/humidity sensor and a passive radiation shield were installed. A new Pyronometer was also installed.

The old Pyronometer and two HMP45C temperature/humidity sensors were shipped to USA for repair and recalibration. A severe corrosion damage beyond repair was found on both circuits of the HMP45C sensors. Power surges were suggested as possible causes. The two HMP45C sensors were declared non-repairable by the Scientific Campbell technicians.  We are advised to replace these with two HC2-S3 sensors. The HMP45Cs are obsolete and will not be manufactured any longer. The Pyronometer was recalibrated and will be shipped back upon payment of the recalibration fee.

 The climate station is now fully operational (less two temperature/humidity sensors). The data looks good. Climate data for November and December was submitted on to the portal. Bi-weekly maintenance of the climate station has been carried out in accordance with the protocol. The maintenance log has also been uploaded onto the portal.

 

Terrestrial vertebrate monitoring:

Images from the camera traps were identified. Data DVDs were sent and received by the TEAM secretariat in Washington. However, we experienced a couple of camera trap failures during this year’s sampling. I suggest that we replace ten of the camera traps to avoid camera failure in the next sampling and eventual data loss.

Protocol Problems: 

Vegetation monitoring:

No problems with protocol implementation.

Climate monitoring:

The climate station was repaired and a few additions were made. This was done in liaison with the TEAM technical personnel. The battery house was modified to allow improved aeration (crucial to ensure battery long life). A new data logger and solar regulator were installed. A new HC2-S3 temperature/humidity sensor and a passive radiation shield were installed. A new Pyronometer was also installed.

The old Pyronometer and two HMP45C temperature/humidity sensors were shipped to USA for repair and recalibration. A severe corrosion damage beyond repair was found on both circuits of the HMP45C sensors. Power surges were suggested as possible causes. The two HMP45C sensors were declared non-repairable by the Scientific Campbell technicians.  We are advised to replace these with two HC2-S3 sensors. The HMP45Cs are obsolete and will not be manufactured any longer. The Pyronometer was recalibrated and will be shipped back upon payment of the recalibration fee.

 The climate station is now fully operational (less with two temperature/humidity sensors). The data looks good. Climate data for November and Decemeber was submitted on to the portal. Bi-weekly maintenance of the climate station has been carried out in accordance with the protocol. The maintenance log has also been uploaded onto the portal.

Terrestrial vertebrate monitoring:

We experienced a couple of camera trap failures during this year’s sampling. I suggest that we replace ten of the camera traps to avoid camera failure in the next sampling and eventual data loss.

Schedule Problems: 

No schedule problems encountered. All activities were implemented on time.

Logistical Problems: 

We experianced a delay to recieve the shipment for our new climate station equipment (data logger and passive radiation shield). TEAM funding was also delayed.

Suggestions: 

No suggestions at the moment.